i think of a number and add 6 to it then multiply the sum by the orignal number and the result is 60.find the number?
Answer by edjones(8007)

Let n be the number.
n(n+6)=60
n^2+6n=60
n^2+6n-60=0
n=sqrt(69)-3
.
Ed